OpenStudy (anonymous):

whats the last digit of 7 to the power of 200

OpenStudy (aaliyahfairgood):

Last digit in 7^200 Here 7 is base and 200 is index to find unit digit, the general method is Base / 4 find remainder, If remainder is 1 then unit digit in the answer is (Base unit digit)^1 is 2 then unit digit in the answer is (Base unit digit)^2 is 3 then unit digit in the answer is (Base unit digit)^3 is 4 then unit digit in the answer is (Base unit digit)^4 Here, 200/4 remainder is zero so unit digit in the answer = 7^4 = 7 × 7 × 7 × 7 = unit digit = 1

OpenStudy (anonymous):

so the answer to that is 1?

OpenStudy (mathmath333):

by binomial theorm \(\large\tt \begin{align} \color{black}{ rem\implies \dfrac{7^{200}}{10}\\~\\ rem\implies \dfrac{49^{100}}{10}\\~\\ rem\implies \dfrac{(50-1)^{100}}{10}\\~\\ rem\implies \dfrac{(5\times 10-1)^{100}}{10}\\~\\ rem\implies \dfrac{(-1)^{100}}{10}\\~\\ rem\implies \dfrac{1}{10}\\~\\ \Large rem\implies 1\\~\\ }\end{align}\)

ganeshie8 (ganeshie8):

nice :)

OpenStudy (zzr0ck3r):

\(7^{200}\mod(10)=\\(7^2)^{100} \mod(10) =\\49^{100}\mod(10) =\\ 9^{100}\mod(10)=\\(9^2)^{50}\mod(10)=\\1^{50}\mod(10)=\\1\mod(10)\)